OpenStudy (lorraine):

how would I write this equation showing all my work y=-2/3x-2 with (-9,4),(-3,0), (6,-6) anyone know

OpenStudy (anonymous):

use the formula..\[\frac{y-y_1}{y_1-y_2}=\frac{x-x_1}{x_1-x_2}\]

OpenStudy (lorraine):

is there any more ways to show me

OpenStudy (anonymous):

putting(x1y1)=(-9,4) and (x2,y2)=(-3,0)

OpenStudy (anonymous):

another way u assume that the equation is ax+by+c=0 put the points and solve for a,b,c

OpenStudy (lorraine):

show me

OpenStudy (anonymous):

put (-9,4) -9a+4b+c=0 put (-3,0) -3a+c=0 put (6,-6) 6a-6b+c=0 now solve these...

OpenStudy (lorraine):

what it says given the equation ye-2/3x-2 find the y coordinates which I did

OpenStudy (lorraine):

how

OpenStudy (lorraine):

y=-2/3

OpenStudy (anonymous):

what is the x value for which u need to find y co-ordinate??

OpenStudy (lorraine):

should I type you the question

OpenStudy (lorraine):

y=-2/3x-2

OpenStudy (anonymous):

\[y=-\frac23x-2\] then??

OpenStudy (lorraine):

yes